OpenStudy (anonymous):

@Nnesha Simplify: (X^2 + 2x + 3) (x^2-2x-2) Help. Fast..

Nnesha (nnesha):

multiply every single term which is in first parentheses by 2nd parentheses

OpenStudy (anonymous):

(x2+2x+3)(x2−2x−2) =(x2+2x+3)(x2+−2x+−2) =(x2)(x2)+(x2)(−2x)+(x2)(−2)+(2x)(x2)+(2x)(−2x)+(2x)(−2)+(3)(x2)+(3)(−2x)+(3)(−2) =x4−2x3−2x2+2x3−4x2−4x+3x2−6x−6 =x4−3x2−10x−6

Nnesha (nnesha):

O_O

OpenStudy (anonymous):

Is that right bc kinda in a hurry here >.<

Nnesha (nnesha):

how about to delete that space?

OpenStudy (anonymous):

lol sorry for the bottom white void and yes thats right!:)

Nnesha (nnesha):

ohh i thought smokeyriver postd that :/

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @SmokeyRiver Is that right bc kinda in a hurry here >.< \(\color{blue}{\text{End of Quote}}\) what right ?

OpenStudy (anonymous):

The answer that he posted^^ And @LightYagami96 Can I post another and tag you in it?

OpenStudy (anonymous):

That is right i am positive and yea sure tag me dont forget the medals:)

OpenStudy (anonymous):

I WONT. I'll fan as well